shishi:

Given a circle with measures of (C, d, and r) and a circle with measures of (C', d', and r'), find r' if C C' = 0.25 and d = 6.0.

darkknight:

what does that mean, c d and r.

darkknight:

like circumference, diameter and radius?

sutralotv12:

I think I see, it is saying that the circumference is 0.25 and the diameter is 6

sutralotv12:

and you have to find the radius

darkknight:

No that can't be right. It isn't that. Because then you can just divide by 2 to find radius. This is something different

darkknight:

Can you send the actual problem, like a screenshot

shishi:

im using a laptop

darkknight:

so?

sutralotv12:

use the snipping too and download it on your computer. I can send you a link on how to download it

darkknight:

yeah

darkknight:

or take a screenshot

shishi:

ok

sutralotv12:

Here is link to how to download snipping tool (it really helps)

sutralotv12:

so is it trying to find the radius using the other two or what?

darkknight:

leave it, when a user doesn't come back to revisit their question when it is confusing then their is no point in wasting time on it

sutralotv12:

ok

sutralotv12:

do you still need this question answered?

shishi:

yes

sutralotv12:

did you get that screenshot? also, if not, could you explain exactly what the problem wants?

shishi:

Given a circle with measures of (C, d, and r) and a circle with measures of (C', d', and r'), find r' if C C' = 0.25 and d = 6.0. A) 6 B) 12 C) 18 D) 24

shishi:

that what it said

darkknight:

no picture?

shishi:

no

shishi:

c/c =0.25

darkknight:

what math is this for may I ask?

shishi:

geometry

darkknight:

there has to be a picture of the circles

darkknight:

what is c, d and r then?

sutralotv12:

I agree

shishi:

these no picture

shishi:

20) Given a circle with measures of (C, d, and r) and a circle with measures of (C', d', and r'), find r' if C/c' = 0.25 and d = 6.0. A) 6 B) 12 C) 18 D) 24

shishi:

that all it said

darkknight:

???

darkknight:

that makes no sense

shishi:

r=raidius

sutralotv12:

is this like 10th grade level geometry because I ain't there yet

sutralotv12:

you question isn't making any sense to me

darkknight:

I passed that, but I don't get it

shishi:

i'm in 11th

darkknight:

incomplete question

darkknight:

@Vocaloid

sutralotv12:

yeah then I ain't there yet and I can't help

shishi:

ok it fine. when if they show me the answer i will post it here

darkknight:

@mhchen

darkknight:

@mhchen do u get this

darkknight:

yeah, like what is the problem

justus:

I agree with the darkknight.

sutralotv12:

It don't make any sense

Vocaloid:

C/C' = d/d' = r/r' due to the similarity between all circles C/C' and d are given, therefore you can solve for d' once you have d', then r' is simply 1/2 d'

sutralotv12:

DAAANGG. We all be up in here thinkin we smort and we can't even solve this problem then Vocaloid come up in here and answer straight away

darkknight:

oh it was asking for thar

darkknight:

thanks voca

sutralotv12:

I feel stupid -_-

shishi:

C C' = d d' 0.25 = 6.0 d' → d' = 24, so r' = 12. The ratio of a circle’s circumference and its diameter is the same for every circle, the ratio is known as π.

shishi:

that the answer

justus:

@shishi if you are done with this post, you can close it with the 'close' button above

Eiwoh2:

There's surely a reason she's earned the Honorary Professor of Mathematics title. \(\color{#0cbb34}{\text{Originally Posted by}}\) @sutralotv12 DAAANGG. We all be up in here thinkin we smort and we can't even solve this problem then Vocaloid come up in here and answer straight away \(\color{#0cbb34}{\text{End of Quote}}\)